prove that sin^2[π÷8+A÷2]-sin^2[π÷8-A÷2]=(1÷√2)sinA

Grade:11

1 Answers

Jayanth Kasaraneni
7 years ago
we know that sin2A-sin2B = sin (A+B)sin(A-B)
π/8 + a/2+π/8 - a/2) sin(π/8 + a/2-π/8+ a/2)sin^2[π÷8+A÷2]-sin^2[π÷8-A÷2]=sin(
π/4) sin(a)=sin(
=(1/\sqrt{2} ) sina